OpenStudy (anonymous):

Write three ratios equal to 5/40

OpenStudy (zehanz):

The ratio is just the same as dividing two numbers, so the same rules also apply. This means if you multiply numerator and denominator with the same number, the ratio doesn't change.

OpenStudy (anonymous):

huh? if I knew how to do that I wouldn't of asked the question.

OpenStudy (zehanz):

Well, that is how I read the question. Ratio 5/40 = 50/400 = 1/8 = 7/ 56 etc. I wouldn't know what else to make of it... Do you have comparable problems, so we can try to understand what is supposed to be done?

OpenStudy (zehanz):

You know, sometimes an obvious question is just that ;)

OpenStudy (anonymous):

Eh I'm not good at math at all even the easiest question. :|

OpenStudy (zehanz):

From now on this kind of question is not a problem anymore, so you've already made progress!

OpenStudy (anonymous):

:| I just guessed on the answer and it was wrong so apparently not.